OpenStudy (nuccioreggie):

The temperature is changing every day by negative 7 over 18 degree Fahrenheit. What will be the change in the temperature after 3 days? 21 over 6 degrees Fahrenheit negative 21 over 6 degrees Fahrenheit negative 7 over 6 degree Fahrenheit 7 over 6 degree Fahrenheit

OpenStudy (anonymous):

when have -7/18 + (-7/18) + (-7/18) = -21/18. simplify -21/18 = -7/6 The answer is C

OpenStudy (anonymous):

@Arock157

OpenStudy (anonymous):

DEEZ NUTZ HOE

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Do you need more help?

OpenStudy (anonymous):

more help with deez nutz